OpenStudy (anonymous):

Please HELP w/ this Word Problem!! Bob drives from A to B at an average speed of 60 mi/h. On the way back, he slowed down to 50 mi/h. The trip was 4.4 hours. What is the distance from A to B? ***please tell me how you can check this problem, i.e. how can you plug in the solution to verify the answer?

OpenStudy (anonymous):

lets call the distance from A to B say \(D\) because that is what you want since \(T=\frac{D}{R}\) that is time is distance divided by rate, you know the time going was \[\frac{D}{60}\] and the time returning was \[\frac{D}{50}\] and the total time is therefore \[\frac{D}{60}+\frac{D}{50}\] which you are told is 4.4 hours, so set \[\frac{D}{60}+\frac{D}{40}=4.4\] and solve for \(D\)

OpenStudy (anonymous):

ooops last line should be \[\frac{D}{60}+\frac{D}{50}=4.4\] a typo on my part

OpenStudy (anonymous):

how would you solve this? I added them to get 100D/3000... is that right?

OpenStudy (anonymous):

nevermind i got it

OpenStudy (anonymous):

How could you check this problem to verify the answer?